A person can row downstream 20 km in 2 hours and upstream 4 km in 2 hours . Find the mans speed of rowing in the still water and speed of the current ? ( sir pls help me . it's urgent THANKS >>>)
Answer by checkley71(8403)

DOWN STREAM 20 KM TAKES 2 HOURS THEREFORE 20/2=10 KM/H
UP STREAM 4 KM TAKES 2 HOURS THEREFORE 4/2=5 KM/H
THEREFORE 10+CURRENT = 5-CURRENT OR 2 CURRENT=5 OR CURRENT =5/2 OR CURRENT=2.5 KM/H